OpenStudy (lexy148):

Use elimination to solve the problem: -4x+2y=22 -x-2y+=18

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \quad ~-4x +2y=22}\) \(\color{#000000 }{ \displaystyle ^{^{\LARGE +}}~~~~-x-2y=18}\) \(\color{#000000 }{ \displaystyle ^\text{__________________________}}\) \(\color{#000000 }{ \displaystyle ~~~~~-5x-0y=40}\)

OpenStudy (solomonzelman):

-5x=40 x=?

OpenStudy (lexy148):

thx for your help i already figured it out

OpenStudy (solomonzelman):

Alright:)